OpenStudy (anonymous):

SHOW ALL OF THE STEP THAT YOU USE TO SOLVE THE PROBLEM 21. Find the distance between points M(6,16) and Z(-1,14) to the nearest tenth.

OpenStudy (anonymous):

Length of a line segment formula: AB √(X2-X1)² + (y2-y1)² MZ = √ (14-16)² + (-1-6)² MZ = √ (-2)² + (-7)² MZ = √ 4 + 49 MZ = √ 53 MZ = 7.280 Round to nearest tenth

OpenStudy (anonymous):

i hope thats helps beautiful

OpenStudy (anonymous):

@samantha1995